# AB-Training: A Communication-Efficient Approach for Distributed Low-Rank Learning

### v1

@@ 6. Conclusion
Our experimental results demonstrate the significant potential of our AB training method to reduce communication overheads by utilizing low-rank representations and independently trained worker subgroups. The consistent 50% reduction in network traffic has the potential to unlock new possibilities within distributed-memory computing environments. This could be particularly impactful for scientific research on machines without high-speed interconnects and for exploring even larger, more complex models. Additionally, the pronounced regularization effects observed at smaller scales offer a promising direction for improving generalization and performance.

### v2

@@ VI Conclusion
Our experimental results demonstrate the significant potential of our AB training method to reduce interconnect traffic by utilizing low-rank representations and independently trained worker subgroups. The consistent 70% reduction in network traffic without an increase in training time nor a decrease in the model’s accuracy has the potential to unlock new possibilities within distributed-memory computing environments. This could be particularly impactful for scientific research on machines without high-speed interconnects and for exploring even larger, more complex models. Additionally, the pronounced regularization effects observed at smaller scales offer a promising direction for improving generalization and performance.